# Some mathematical issues regarding a new approach towards quantum foundation

### v2

@@ 1 Introduction
In Helland (2024a) it was assumed that a multivariate representation $U(\cdot)$ with certain properties existed. In the present paper, we show how such a simple representation can be constructed. This assumption is simply not necessary.
From this, the weakest possible version of my main theorem is given as Theorem 1 below. The conclusion of the theorem is that every accessible variable has a symmetric operator in $\mathcal{H}=L^{2}(\Omega_{\psi},\nu)$ attached to it, where $\psi=(\theta,\eta)$, and $\nu$ is an invariant measure on $\Omega_{\psi}$ induced by the invariant measure $\mu$ on $\Omega_{\theta}$. This is the starting point for much of the Hilbert space apparatus. Apart from the above symmetry assumption, the essential assumption is only the existence of two complementary theoretical variables in the given context.
Note that there are no macroscopic assumptions here. Thus, this derivation also gives a foundation for what Khrennikov (2010, 2023) calls quantum-like models. These models have links to several scientific disciplines. The link to quantum decision theory will be discussed elsewhere. Links to relativity theory and quantum field theory are discussed in Helland (2023c) and in Helland and Parthasarathy (2024).
I will concentrate on the derivation of the Hilbert space apparatus and related derivations in this article. Assumptions that lead to the Born rule for probabilities are discussed, and the derivation is proved, in Helland (2024b). Derivation of the Schrödinger equation from a few postulates has been given for instance by Klein (2010).
Finally, it is a basic setting in that the theoretical variables are attached to an observer or to a group of communicating observers. A natural assumption in addition is that they are connected to the mind(s) of this/these observer(s). This leads to an epistemic interpretation of quantum theory, an interpretation that contains QBism (see for instance Fuchs et al., 2013) as a special case. Quantum theory is seen as a theory of our knowledge about the real world, not directly about the real world.
@@ 2 The main theorems
I repeat that my main notion is that of theoretical variables, which can be almost anything. The theoretical variables can be accessible or inaccessible. From a mathematical point of view, I only assume that if $\lambda$ is a theoretical variable and $\theta$ is a fixed function of $\lambda$, then $\theta$ is a theoretical variable. And if $\lambda$ is accessible, then $\theta$ is accessible.
@@
In simple physical examples, such a $\phi$ can easily be constructed. As a general statement covering all possible situations, Postulate 1 can also be given a religious interpretation, see Helland (2022d, 2023d).
@@
This is a partial ordering among all theoretical variables, and also a partial ordering among the accessible ones. Note that $\phi$ from Postulate 1 is an upper bound in the accessible case. We will say that $\theta$ is a maximal accessible variable if it is maximal with respect to this partial ordering. By Zorn’s Lemma, which is equivalent to the Axiom of Choice, maximal variables always exist. For those who do not believe in Zorn’s Lemma, we add an additional postulate.
@@
It should also be mentioned that, in addition to the Postulates of this article and of Helland (2024b,c), a final postulate is needed to compute probabilities of independent events. A version of such a postulate is
Postulate 5
If the probability of an event $E_{1}$ is computed by a probability amplitude $z_{1}$ from the Born rule in the Hilbert space $\mathcal{H}_{1}$, the probability of an event $E_{2}$ is computed by a probability amplitude $z_{2}$ from the Born rule in the Hilbert space $\mathcal{H}_{2}$, and these two events are independent, then the probability of the event $E_{1}\cap E_{2}$ can be computed from the probability amplitude $z_{1}z_{2}$, associated with the Hilbert space $\mathcal{H}_{1}\otimes\mathcal{H}_{2}$.
This postulate can be motivated by its relation to classical probability theory: If $P(E_{1})=|z_{1}|^{2}$ and $P(E_{2})=|z_{2}|^{2}$, then
$P(E_{1}\cap E_{2})=P(E_{1})P(E_{2})=|z_{1}|^{2}|z_{2}|^{2}=|z_{1}z_{2}]^{2}.$
In Theorem 1 it was concluded that the relevant operators were symmetric. This is a simple property: $\langle u|Av\rangle=\langle Au|v\rangle$ for all $|u\rangle,|v\rangle$ in the domain of $A$. To be precise, we need operators corresponding to the two maximal accessible variables $\theta$ and $\eta$ are self-adjoint. Then the spectral theorem is valid (Hall, 2013), and it can be used to define operators corresponding to other accessible variables.
@@
If Postulate 6 holds, then $A^{\theta}$ is self-adjoint. A corresponding condition holds for $A^{\eta}$.

### v3

@@ 1 Introduction
From a mathematical point of view, the notion of a theoretical variable will be considered as primitive. The only requirement is that if $\theta$ is a theoretical variable and $\xi$ is a function of $\theta$, then $\xi$ is a theoretical variable. Similar, the notion of an accessible variable is primitive. Again, it is only required that if $\theta$ is accessible and $\xi$ is a function of $\theta$, then $\xi$ is accessible.
Given these primitive notions, the theory may be implemented in different ways. In this article, I will concentrate on the case where the variables are physical variables like positions, momenta or spin components. A physical variable is said to be accessible if it in principle can be measured with arbitrary accuracy.
Other implementations are possible. In Helland (2023e) the theoretical variables are decision variables. In Helland (2025) they are statistical parameters. In both cases, a fairly rich applied theory can be constructed from the mathematical theory considered here.
@@
In Helland (2024a) it was assumed that a multivariate representation $U(\cdot)$ of $G$ with certain properties existed. In the present paper, we show how such a simple representation can be constructed. This assumption is simply not necessary.
From this, the weakest possible version of my main theorem is given as Theorem 1 below. The conclusion of the theorem is that every accessible variable has a symmetric operator in $\mathcal{H}=L^{2}(\Omega_{\psi},\nu)$ attached to it, where $\psi=(\theta,\eta)$, and $\nu$ is an invariant measure on $\Omega_{\psi}$ induced by the invariant measure $\mu$ on $\Omega_{\theta}$. (Strictly speaking, for this conclusion we need technical assumptions such that the spectral theorem is valid for symmetric operators, not only for self-adjoint operators.) This is the starting point for much of the Hilbert space apparatus. Apart from the above symmetry assumption, the essential assumption is only the existence of two complementary theoretical variables in the given context.
Note that there are no microscopic assumptions here. Thus, this derivation also gives a foundation for what Khrennikov (2010, 2023) calls quantum-like models. These models have links to several scientific disciplines. The link to quantum decision theory will be discussed elsewhere. Links to relativity theory and quantum field theory are discussed in Helland (2023c) and in Helland and Parthasarathy (2024).
I will concentrate on the derivation of the Hilbert space apparatus and related derivations in this article. Assumptions that lead to the Born rule for probabilities are discussed, and the derivation is proved, in Helland (2024b). The assumptions are also discussed in Section 3 below. Derivation of the Schrödinger equation from a few postulates has been given for instance by Klein (2010).
Finally, it is a basic setting in that the theoretical variables are attached to an observer or to a group of communicating observers. A natural assumption in addition is that they are connected to the mind(s) of this/these observer(s). This leads to a general epistemic interpretation of quantum theory, an interpretation that contains QBism (see for instance Fuchs et al., 2013) as a special case. Quantum theory is seen as a theory of our knowledge about the real world, not directly about the real world. This is discussed in Section 4.
@@ 2 The main theorems
I repeat that my main notion is that of theoretical variables, which can be almost anything. The theoretical variables can be accessible or inaccessible.
@@
In simple physical examples, such a $\phi$ can easily be constructed. As a general statement covering all possible situations, Postulate 1 can also be given a religious interpretation, see Helland (2022d, 2023d). This will be further discussed elsewhere.
@@
This is a partial ordering among all theoretical variables, and also a partial ordering among the accessible ones. Note that $\phi$ from Postulate 1 is an upper bound in the accessible case. (Again, one must make precise what is meant by a function. For instance, one can let $\Omega_{\phi}$ and $\Omega_{\theta}$ be topological spaces, and concentrate on functions that are Borel-measurable.) We will say that $\theta$ is a maximal accessible variable if it is maximal with respect to this partial ordering. By Zorn’s Lemma, which is equivalent to the Axiom of Choice, maximal variables always exist. For those who do not believe in Zorn’s Lemma, we add an additional postulate.
@@
In Theorem 1 it was concluded that the relevant operators were symmetric. This is a simple property: $\langle u|Av\rangle=\langle Au|v\rangle$ for all $|u\rangle,|v\rangle$ in the domain of $A$. To use the spectral theorem in general, we need operators corresponding to the two maximal accessible variables $\theta$ and $\eta$ are self-adjoint. Then the spectral theorem is valid (Hall, 2013), and it can be used to define operators corresponding to other accessible variables.
@@
If Postulate 5 holds, then $A^{\theta}$ is self-adjoint. A corresponding condition holds for $A^{\eta}$.
@@
3 Quantum probabilities
For the derivation of Born’s formula, two more postulates are needed.
Postulate 6
The generalized likelihood principle from statistics holds.
The basis for nearly all statistical inference is a statistical model, a model for the data $z$ in terms of the total parameter $\theta$. This is expressed by a probability function $p(z|\theta)$, in the discrete case a point probability, and in the continuous case a probablity density. The likelihood of the data is defined as $L(\theta|z)=p(z|\theta)$, the probability function seen as a function of the parameter.
The following principle can be taken as the basis for very much statsitical inference, and in my theory it is also used as a motivation behind Born’s formula.
The generalized likelihood principle
Consider two experiments in the same context $\tau$, and assume that $\theta$ is the same full parameter in both experiments. Suppose that two observations $z_{1}$ and $z_{2}$ have proportional likelihoods in the two experiments, where the proportionality constant $c$ is independent of $\theta$. Then, these two observations produce the same experimental evidence on $\theta$ in this context.
The term “experimental evidence” is here left undefined and can be specified in any desirable direction. In some statistical textbooks and articles, the likelihood principle is formulated without specifying a context. As discussed in Helland (2021, 2024b), specifying $\tau$ implies that the principle becomes much less controversial. One may also assume that the two contexts are similar, i.e., one-to-one functions of each other.
The second necessary postulate is
Postulate 7
Consider a physical context $\tau$ observed by an observer $B$ whose decisions are influenced by a superior actor $D$. Assume that $D$’s probabilities are taken as experimental evidence and that $D$ is seen by $B$ to be rational in agreement with the Dutch Book Principle.
The Dutch Book Principle
No choice of payoffs in a series of bets shall lead to a sure loss for the bettor.
From these two postulates, assuming perfect experiments in the sense that no experimental noise is assumed, and using a version of Gleason’s theorem given by Busch (2003), the following versions of Born’s formula is proved.
Consider two experiments with maximal accessible discrete parameters $\theta^{a}$ and $\theta^{b}$, having the same number of values. By a corollary to Theorem 1 (Helland, 2024a), there are operators $A^{a}$ and $A^{b}$ in a Hilbert space $\mathcal{H}$ associated with these two parameters, the eigenvalues of the operators are the possible values of the parameters, and the corresponding normed eigenvectors are simple. Let $|a;k\rangle$ correspond to the event $\theta^{a}=u_{k}$ and $|b;j\rangle$ correspond to the event $\theta^{b}=v_{j}$.
Theorem 4
The conditional probabilities are given by
$P(\theta^{b}=v_{j}|\theta^{a}=u_{k})=|\langle a;k|b:j\rangle|^{2}.$
Thus, the probabilities are given by the squared norms of the probablity amplitudes $c=\langle a;k|b:j\rangle$.
From this, more general variants of the Born formula can be proved, for instance the following, valid both for discrete and continuous variables, and taking as a point of departure a density operator $\rho^{a}$ computed from an assumed probability distribution over $\theta^{a}$:
Theorem 5
$E(\theta^{b}|\rho^{a})=\mathrm{trace}(\rho^{a}A^{b}).$
A referee has asked the following question: Why did we get precisely the complex Hilbert space formalism, and not the generalized probability theory, in the spirit of Khrennikov et al. (2025)? The generalized probability theory is described in detail in Barrett (2007) and Müller (2021), but we do not need the definitions here.
The answer to why quantum probablities in my theory are given by the Born formula, and not by any more generalized formula, is given by the proofs of the results above, and in particuar by a consequence of the generalized likelihood principle. In the discrete case, one can define the likelihood effects
$F^{b}(\bm{v};z^{b},\tau)=\sum_{j}p(z^{b}|\tau,\theta^{b}=v_{j})|b;j\rangle\langle b;j|,$ (3)
where again $|b;j\rangle$ corresponds to the event $\theta^{b}=v_{j}$. These likelihood effects are crucial in the proof of Theorem 4. They are closely related to the positive operator-valued measures (POVM) of traditional quantum theory, where for discrete data the measure of the event $z^{b}=z$ is given by $\sum_{j}p(z^{b}=z|\tau,\theta^{b}=v_{j})|b;j\rangle\langle b;j|$.
Thus, my approach towards probabilities lead to traditional quantum probabilities.
It should also be mentioned that, in addition to the postulates of this article, a final postulate is needed to compute probabilities of independent events. A version of such a postulate is
Postulate 8
If the probability of an event $E_{1}$ is computed by a probability amplitude $c_{1}$ from the Born rule in the Hilbert space $\mathcal{H}_{1}$, the probability of an event $E_{2}$ is computed by a probability amplitude $c_{2}$ from the Born rule in the Hilbert space $\mathcal{H}_{2}$, and these two events are independent, then the probability of the event $E_{1}\cap E_{2}$ can be computed from the probability amplitude $c_{1}c_{2}$, associated with the Hilbert space $\mathcal{H}_{1}\otimes\mathcal{H}_{2}$.
This postulate can be motivated by its relation to classical probability theory: If $P(E_{1})=|c_{1}|^{2}$ and $P(E_{2})=|c_{2}|^{2}$, then
$P(E_{1}\cap E_{2})=P(E_{1})P(E_{2})=|c_{1}|^{2}|c_{2}|^{2}=|c_{1}c_{2}]^{2}.$
4 The epistemic interpretation and Ozawa’s intersubjectivity theorem
The interpretation of quantum theory is still a theme of intensive debate and some confusion. In a recent poll among a group of physicists (Jodlicska et al. 2025), 21% supported an epistemic interpretation, while 24% supported an ontic interpretation. 15% supported a mix of the epistemic and the ontic interpretation, while 41% said that they supported an ensemble interpretation. In Helland (2019, 2021), I gave several arguments for the epistemic viewpoint. These arguments were largely of non-mathematical nature, and so, the main discussion there is outside the scope of the present article. But one special issue in that connection contains some important mathematics.
In Ozawa (2019), the author proved the following: Assume two spacelike measurements $M_{1}$ and $M_{2}$ at times $\tau_{1}$ and $\tau_{2}$ of the same observable, associated with the operator $A$. Let $U(t)$ be the time evolution operator of the total system, i.e., physical system plus environment. Define $A(0)=A\otimes I$, $M_{1}(t)=U(t)^{\dagger}(I\otimes M_{1})U(t)$ and $M_{2}(t)=U(t)^{\dagger}(I\otimes M_{2})U(t)$. Assume that for all $x$
$P(M_{1}(\tau_{1})=x)=P(M_{2}(\tau_{2})=x)=P(A(0)=x).$ (4)
Then
$P(M_{1}(\tau_{1})=x,M_{2}(\tau_{2})=y)=0$ (5)
for $x\neq y$.
In other words, under the assumptions of space separabiliy and a weak reproducability condition, the two obsevers are forced to get equal measurements.
Khrennikov (2024a) recently argued that the QBism’s fundamental statement that ‘the measurement of an observable is personal’ is in contraversion with Ozawa’s theorem. His starting point is the following citation from Fuchs and Schack (2014): ‘The fundamental primitive of QBism is the concept of experience. According to QBism, quantum measurement is a theory that any agent can use to eveluate her expectations for the content of their experience.’
Following Khrennikov, Ozawi’s theorem implies that, for accurate local observations, measurement’s outcome is intersubjective which is a strong objection to QBism. There is nothing concernng personal experiences when different observers by necessity obtain the same measurement result.
This should be contrasted to my own interpretation, which can be seen as a consequence of the mathemathical theory sketched in Sections 2 and 3 above. This mathematical theory can, in a physical context be interpreted in two ways: First, the mathematics, including the Born formula, can be seen as what is given by the observations of a single observer. It should be emphasized that the probabilities in Born’s formula can be seen as probabilities calculated by the hypothetical superior actor $D$. Next, the mathematics can be interpreted as the joint observations as seen by a group of communicating observers.
Note that, in Ozawa’s theorem, the two observers are specelike separated, and do not communicate about their observations. However, spacelike separated observers may later meet and share their experiences. Then it is perfectly allowable that they find that their marurement values are equal.
In conclusion, their seems to be no contraversion between my general epistemic theory and Ozawa’s intersubjectivity theorem.
It is interesting that Khrennikov (2024b) argues that Rovelli’s well-knows relational quantum mechanics can be specified in a way where there is no contrdiction to Ozawa’s theorem. Relational quantum mechanics states that every statement about a system should be related to another system. This ‘other system’ may well be an observer or a communicating group of observers, which provides a link to my interpretation.